Write 18765^12 as a mathematical expression without using the number itself and without using scientific notation
Answer by swincher4391(1107)

Well 18765^12 = (18765^2)^6 = (352125225)^6.
The number gets pretty out of hand if we raise 18765 to the 3rd power.